OpenStudy (anonymous):

Find all values of x for which the function y=(2(8-5x))/(x+1)

OpenStudy (anonymous):

i think question is incomplete !

OpenStudy (anonymous):

is differentiable.

OpenStudy (anonymous):

y=([\sqrt{8-5x}\])/(x+1)...it's a square root also..not a two

OpenStudy (anonymous):

\[f(x)=\frac{\sqrt{8-5x}}{x+1}\]

OpenStudy (anonymous):

yes, that's what it looks like!

OpenStudy (anonymous):

ok u must check for continuity first

OpenStudy (anonymous):

when I graphed it, it looked like 1/x graph

OpenStudy (anonymous):

yeah its something like that..maybe

OpenStudy (anonymous):

so how do I check for continuity?

OpenStudy (anonymous):

ok u have some restrictions here.. division by zero is not allowed under radical cant be negative

OpenStudy (anonymous):

actually u must find domain of ur function

OpenStudy (anonymous):

domain is x=-1

OpenStudy (anonymous):

\[x\neq -1\]

OpenStudy (anonymous):

ohyeah you're right

OpenStudy (anonymous):

and what about radical?

OpenStudy (anonymous):

what's radical?

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

x\[\neq5/8\]

OpenStudy (anonymous):

er 8/5

OpenStudy (anonymous):

when u have something like \[\sqrt{x}\] u must have \[x\ge0\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what about the expression under square root sign?

OpenStudy (anonymous):

\[x \neq 8/5\]

OpenStudy (anonymous):

emm i mean u must have \(8-5x\ge0\) and this leads to\[x\le\frac{8}{5}\]make sense?

OpenStudy (anonymous):

oh yes sorry i meant to do that sign instead of the nonequal sign

OpenStudy (anonymous):

it does make sense

OpenStudy (anonymous):

good

OpenStudy (anonymous):

ok so the domain of function becomes\[x\le \frac{8}{5} \ \ \text{and} \ \ x\neq-1\]

OpenStudy (anonymous):

or u can write it like this\[(-\infty,-1)\cup (-1,\frac{8}{5}]\]

OpenStudy (anonymous):

got it. so is that where it's differentiable?

OpenStudy (anonymous):

yes thats it ... and just one more point function is not continous at end points \(x=-1\) and \(x=\frac{8}{5}\) so it'll not be differentiable at that points...

OpenStudy (anonymous):

ok thank you so much! i appreciate it :)

OpenStudy (anonymous):

very welcome :)